Calculate the Hall coefficient of sodium 2 points based on free electron model. Sodium has bcc structure and the side of the cube is 4.28 Å.
1
Expert's answer
2021-09-22T09:23:28-0400

The expression for Hall coefficient:


"R_H=\frac 1{ne}."

The concentration density is


"n=\frac{\rho a^3N_A}{\Mu},"

where

"\rho" - density,

a - cell parameter,

NA - Avogadro constant,

M - molar mass.


"R_H=\frac{\Mu}{a^3e\rho N_A}=3.13·10^{18}\text{m}^3\text{/C}."